OpenStudy (anonymous):

solve for x: -6(3x-8)=-6

OpenStudy (asapbleh):

-18x+48=-6 -18x=-54 x=3

OpenStudy (asapbleh):

Check: -18(3)+48=-6 It is correct

OpenStudy (anonymous):

thanxs @asapbleh

OpenStudy (anonymous):

no its not @asapbleh